OpenStudy (anonymous):

Convert 310/20 to a mixed number. 15 1/2 15 3/4 14 1/2 14 3/4

OpenStudy (agent47):

310/20 = (300+10)/20 = 15 1/2

OpenStudy (anonymous):

310/20=15 1/2

OpenStudy (anonymous):

Thanks @Agent47 and @wangkan07 :)